Christopher is 27 years younger than Vanessa. For the last two years, Vanessa and Christopher have been friends. Six years ago, Vanessa was 4 times as old as Christopher. How old is Christopher now?
Answer by ankor@dixie-net.com(22740)

Christopher is 27 years younger than Vanessa.
c = v - 27
or
v = c + 27
:
Six years ago, Vanessa was 4 times as old as Christopher.
v - 6 = 4(c - 6)
v - 6 = 4c - 24
v = 4c - 24 + 6
v = 4c - 18
:
How old is Christopher now?
Replace v with (c+27) from the first statement
c + 27 = 4c - 18
c - 4c = -18 - 27
-3c = -45
c = -45/-3
c = +15 yrs is Christopher's age
:
:
:
Confirm this in the 2nd statent
v = 15 + 27 = 42
42 - 6 = 4(15-6)
36 = 4(9)